OpenStudy (anonymous):

Given f(x)=6x and g(x) = x^2-4x Find (f+g)(x)

OpenStudy (kira_yamato):

(f + g)(x) = f(x) + g(x)

OpenStudy (anonymous):

just add them together?

OpenStudy (anonymous):

(f+g)(x) = x^2 +2x

OpenStudy (anonymous):

I think my answer would be x^2 +2x ?

OpenStudy (ddcamp):

Yup, x²+2x is right.

OpenStudy (anonymous):

Ty